### v1

# Rydberg Atomic Receiver: Next Frontier of Wireless Communications

Abstract
The advancement of Rydberg Atomic REceiver (RARE) is driving a paradigm shift in electromagnetic (EM) wave measurement. RAREs utilize the electron transition phenomenon of highly-excited atoms to interact with EM waves, thereby enabling wireless signal detection. Operating at the quantum scale, such new receivers have the potential to breakthrough the sensitivity limit of classical receivers, sparking a revolution in physical-layer wireless communications. The objective of this paper is to offer insights into RARE-aided communication systems. We first provide a comprehensive introduction to the fundamental principles of RAREs. Then, a thorough comparison between RAREs and classical receivers is conducted in terms of the antenna size, sensitivity, coverage, and bandwidth. Subsequently, we overview the state-of-the-art design in RARE-aided wireless communications, exploring the latest progresses in frequency-division multiplexing, multiple-input-multiple-output, wireless sensing, and quantum many-body techniques. Finally, we highlight several wireless-communication related open problems as important research directions.
Index Terms:
Rydberg atomic receivers, wireless communications, quantum sensing.
I Introduction
The six-generation (6G) communication networks will leverage all available resources to support the explosive growth of global mobile services [1]. To achieve seamless coverage, 6G will expand beyond terrestrial networks in 1G–5G to space–air–ground integrated networks (SAGINs) [1]. Furthermore, the full spectrum, including sub-6GHz, millimeter-wave (mmWave), and terahertz (THz) bands, will be explored for attaining high-rate and multi-modal connections. These ambitious long-range and full-frequency communication visions require the next-generation receiver to exhibit exceptional sensitivity to electromagnetic (EM) waves across a wide spectrum.
Since Hertz validated the existence of EM waves, classical receivers detect wireless signals primarily by using an antenna to transform incident EM waves into informative electric currents [2]. This mechanism, however, presents insurmountable limitations to physical-layer (PHY) wireless communications. Specifically, the front-end circuits of a receiver, such as amplifiers and mixers, introduce thermal noise into data currents. The noise power, -174 dBm/Hz, restricts the receiver’s sensitivity limit. In addition, an ideal dipole antenna typically has a half-wavelength size, which constrains its operating bandwidth to a narrow range around the center carrier frequency [1]. To detect signals across the sub-6GHz, mmWave, and THz bands, the receiver has to equip an array of bulky antennas and font-end circuits, each dedicated to a specific band. Consequently, these challenges constrain the ultimate performance of wireless communications.
Recently, at the intersection of quantum sensing and wireless communication, the Rydberg Atomic REceiver (RARE) has emerged as a radical solution to overcome these limitations [3]. Rydberg atoms are highly-excited atoms with outer electrons far from the nucleus. Due to the large electric dipole moments, Rydberg atoms can strongly interact with incident EM waves, triggering electron transitions between resonate energy levels. As a result, RAREs can accurately detect wireless signals by monitoring changes on the atoms’ quantum states resulting from electron transitions [4]. Representing a paradigm shift in receiver architecture to embrace quantum sensing, RAREs are gaining increasing attentions due to their quantum advantages. For instance, a RARE is immune to thermal noise as it involves no front-end circuits. Hence, its sensitivity limit can far exceed those of classical counterparts [5]. Moreover, a single RARE is able to resonate with a large number of frequencies ranging from MHz to THz band because of the abundant energy levels in Rydberg atoms [6]. These quantum advantages make RAREs a promising technology to revolutionize PHY communications in diverse applications, such as weak signal detection, coverage enhancement, power saving, and full-frequency sensing.
The objective of this paper is to offer a beginner’s guide to the next-generation wireless communications aided by RAREs. To this end, we first comprehensively introduce the fundamental features of Rydberg atoms and the key modules of a RARE. Then, the performance of RAREs is compared with classical receivers in antenna size, sensitivity, coverage, and bandwidth. Subsequently, we overview the state-of-the-art RARE techniques for wireless communications. The latest progresses in four important directions are covered: frequency-division-multiplexing (FDM), multiple-input-multiple-output (MIMO), wireless sensing, and quantum many-body techniques. Last, we highlight several wireless-communication related open problems as important future directions to conduct research on RAREs.
II Fundamentals of Rydberg Atomic Receivers
Figure 1: Framework of RARE-enabled wireless communications. The “Microscopic view” adopts the structure of Rb atom as an example.
We commence by introducing the fundamentals of RAREs, ranging from the preliminary of Rydberg atom to the key modules of a RARE.
II-A Preliminary of Rydberg Atom
II-A1 Definition of Rydberg atom
An atom consists of a positively charged nucleus orbited by negatively charged electrons. According to the Bohr model, electrons in these orbits have unique quantum states and quantized energy levels. Each orbit is labeled by three quantum numbers $n\ell_{j}$, (e.g., $n\ell_{j}=5S_{1/2}$), including a principal quantum number $(n=1,2,3,\cdots)$, an azimuthal quantum number $(\ell=S,P,D,\cdots)$, and a total angular momentum quantum number $(j=\frac{1}{2},\frac{3}{2},\frac{5}{2},\cdots)$. While the latter two numbers, $\ell$ and $j$, determine the angular momentum and orbital shape, the principal quantum number, $n$, affects the orbital radius and energy. For a higher value of $n$, the electron is farther from the nucleus and has higher energy (see Fig. 1). Then, Rydberg atom is defined as a highly-excited atom with one or more electrons that have a large principal quantum number, i.e., $n>20$ [4].
II-A2 Electron transition for preparing Rydberg atom
Electrons can transit between energy levels by either absorbing or emitting photons from EM waves (or lasers) [5]. This electron-transition process occurs when the photon’s frequency resonates (or nearly resonates) with the transition frequency between energy levels, which is defined as the ratio between the energy-level spacing and Plank constant. The process facilitates the creation of Rydberg atoms via the excitation of electrons to high energy levels. This is typically accomplished by counter-propagating two lasers, namely a probe laser and a coupling laser, through a vapor cell filled with Alkali atoms (e.g., Rb and Cs), as shown in Fig. 1 [3]. The probe laser excites the outermost electrons from the ground state to a lowly-excited energy level, (e.g., $5S_{1/2}\rightarrow 5P_{3/2}$). The coupling beam further triggers the electron transition to a Rydberg state, (e.g. $5P_{3/2}\rightarrow 47S_{5/2}$), thereby creating Rydberg atoms. Customarily, the three involved energy levels, (e.g., $5S_{1/2}$, $5P_{3/2}$, $47S_{5/2}$), are denoted as |$\ket{1}$, |$\ket{2}$, and |$\ket{3}$ [4].
II-A3 Properties of Rydberg atom
Given the high principal quantum number $n$, Rydberg atoms exhibit useful properties for sensing EM waves. First, the radius of an Rydberg atom is large due to its dependence on $n^{2}$ [4]. For instance, the atomic radius grows from $150a_{0}$ to $10^{4}a_{0}$ as $n$ increases from $10$ to $100$, with $a_{0}=5.29\times 10^{-11}{\rm m}$ being the Bohr radius. The substantial increase in radius endows the Rydberg atom with a large electric dipole moment, thus making it very sensitive to external EM wave. Second, the transition frequency between adjacent Rydberg states is abundant as the energy level spacing scales as $n^{-3}$ [3]. For low energy levels ($n<20$), the transition frequency can reach the visible light frequency ($>100$ THz). While for Rydberg states ($n>20$), the transition frequencies fall within the range of MHz to THz [5], exactly covering all the typical frequency bands used in PHY communications.
II-B Key Modules of RAREs
The extreme sensitivity of Rydberg atoms to EM waves ranging from MHz to THz motivates the development of RAREs for EM-wave measurement. A RARE typically comprises three key modules: electron-transition based antennas, Electromagnetically-Induced-Transparency (EIT) based quantum measurement, and receive signal processing.
II-B1 Electron-transition based antennas
[…]
II-B2 EIT based quantum measurement
Figure 2: EIT for measuring AM/FM/PM signals. For AM and PM detection, the detuning $\delta$ is zero, while for FM detection, the detuning $\delta$ is nonzero.
To retrieve the wireless signal, it is essential to accurately readout the value of Rabi frequency. The standard measurement scheme is the EIT spectroscopy, developed based on the following two observations [9].
  • The $(1,2)$-th element of the final quantum state, $\rho_{12}$, determines the rate at which Rydberg atoms absorb photons from the probe laser. The absorption loss of the probe-laser scales exponentially with $\rho_{12}$ [7]. Thereby, it is feasible to deduce the value of $\rho_{12}$ by measuring the output power of the probe laser using a photodetector (PD), as depicted in Fig. 1.
  • The quantum state, $\rho_{12}$, has a one-to-one mapping with the Rabi frequency. Specifically, $\rho_{12}$ displays two near-zero points at two specific probe-laser frequencies situated around the transition frequency $f_{\ket{1}\rightarrow\ket{2}}$. The interval of these two frequencies is shown to be linearly proportional to the Rabi frequency $\Omega(t)$ [3].
These two observations suggest that by sweeping the probe-laser frequency around $f_{\ket{1}\rightarrow\ket{2}}$, two peaks can appear in the spectrum of the output probe laser, as depicted in Fig. 2. This allows the Rabi frequency to be measured from the peak interval, which is the principle of EIT-based quantum measurement.
II-B3 Receive signal processing
After obtaining the value of Rabi frequency, the information modulated on ${\mathbf{E}}(t)$ can be recovered via customized signal processing implemented on traditional digital processors. As in Fig. 2, both amplitude-modulated (AM) and frequency-modulated (FM) symbols can be illustrated directly using the maximum likelihood method due to the dependence of Rabi frequency on $\left|{\bm{\mu}^{T}{\mathbf{E}}(t)}\right|$ and $\delta$. For the detection of phase-modulated (PM) symbols, a EIT-heterodyne scheme was developed in [7], which originates from the holographic-phase sensing technique in optics imaging. It introduces an auxiliary source to produce a known reference signal, ${\mathbf{E}}_{r}(t)$, which interferes with the unknown data signal, ${\mathbf{E}}(t)$, over the air. Their radio frequencies can be either identical or a few kHz apart. Consequently, the measured Rabi frequency changes from $|\bm{\mu}^{T}{\mathbf{E}}(t)|$ to $|\bm{\mu}^{T}{\mathbf{E}}(t)+\bm{\mu}^{T}{\mathbf{E}}_{r}(t)|$ (see the bottom-subfigure in Fig. 2). Accordingly, the information modulated on the global phase of ${\mathbf{E}}(t)$ can be extracted from the relative phase between ${\mathbf{E}}(t)$ and ${\mathbf{E}}_{r}(t)$ using spectrum analysis methods or phase retrieval algorithms [8].
So far, we have discussed the mechanism of detecting AM/FM/PM wireless signals using RAREs. Recent studies have also reported the detection of multi-frequency signals [10], multi-user signals [8], and target vibrations [5], all based on the principles we have covered and will be elaborated in Section IV.
III Rydberg Atomic Receiver vs. Classical Receiver
The RARE technique represents a paradigm shift in wireless signal detection from using electronic circuits to embracing quantum sensing. This gives rise to a range of unique features for RARE-enabled PHY communications, which are introduced in this section.
III-A Antenna Size
According to the Chu limit, a lossless dipole antenna in a classical receiver (CR) has a size of half-wavelength, $\frac{\lambda}{2}$. This wavelength-dependent size leads to bulky receivers at low frequencies, e.g., exceeding 10 cm when $f<1$ GHz. In contrast, the vapor cells (i.e., antennas of RAREs) have wavelength-independent sizes, e.g., typically fixed as $1$ cm in diameter [7]. Consequently, RAREs can have electrically small antennas, offering significant advantages especially for low-frequency applications.
III-B Sensitivity Limit
The sensitivity of a receiver, $E_{\min}$, defined as the minimum detectable field strength per square root of averaging time (unit: $\rm V/m/\sqrt{Hz}$), is fundamentally limited by the internal noise generated by the receiver itself. For an ideal $\frac{\lambda}{2}$ dipole antenna of a CR, its internal noise stems from the thermal noise generated by front-end circuits, such as amplifiers and mixers. The corresponding sensitivity limit is expressed as $E_{\min}=\sqrt{\frac{P_{N}Z_{0}}{A_{e}}}$, where $P_{N}=-174{\rm dBm}/{\rm Hz}$ denotes the room-temperature thermal-noise power, $Z_{0}$ the free-space impedance, and $A_{e}$ the effective area. On the contrary, since a RARE contains no front-end circuits, its thermal noise is significantly weaker than a CR [5]. Its major internal noise is the quantum shot noise (QSN) resulting from the randomness when observing quantum states [2]. According to the quantum sensing theory, the QSN power is inversely proportional to the electric dipole moment, $\bm{\mu}$, the number of participating atoms, $N_{a}$, and the coherence time, $T_{r}$ of the EIT process. The theoretical sensitivity limit of a RARE, also called Standard Quantum Limit (SQL), is calculated as $E_{\min}=\frac{\hbar}{|\bm{\mu}|\sqrt{N_{a}T_{r}}}$ [2].
Figure 3: Sensitivity comparison between RAREs and CRs.
Fig. 3 compares the sensitivity of RAREs with CRs. Representative values $T_{r}=225{\rm\mu s}$ and $N_{a}=5\times 10^{5}$ as in [3] are utilized. The dipole moment, $\bm{\mu}$, is selected from the electron transition $nD_{5/2}\leftrightarrow(n+1)P_{3/2}$ that resonates with each frequency. For a fair comparison, we also plot the sensitivity limit of a 1-cm dipole antenna calculated in [2] as it has a similar size with a vapor cell. Fig. 3 illustrates that the SQL can surpass the sensitivity limit of $\frac{\lambda}{2}$ dipole antennas by orders of magnitude. This is attributed to the large electric dipole moment of Rydberg atoms. Moreover, since the early exploitation of RARE in [9] (Oklahoma Univ.@2012), the RARE’s sensitivity has been improved by $10^{4}$ times. Recently developed RAREs have outperformed similarly sized dipole antennas and achieved sensitivities comparable to ideal $\frac{\lambda}{2}$ counterparts. The state-of-the-art design in [11] (Warsaw Univ.@2024) pushes the sensitivity down to $0.4{\rm\mu V/m/\sqrt{Hz}}$ at 13.9 GHz by using six-wave-mixing techniques to remove background noises. It is anticipated that in the near future, RAREs will surpass the ideal $\frac{\lambda}{2}$ dipole antenna and approach SQL, with multi-wave-mixing emerging as a promising approach for performance improvement.
III-C Long Coverage Range or Low Transmission Power
Given the extreme sensitivity promised by SQL, RAREs can theoretically detect much weaker signals than CRs, making it possible to significantly expand communication coverage. To see it more clearly, we plot Fig. 4 to illustrate the feasible communication distance, $d$, with respect to (w.r.t) the transmit power, $P_{t}$, for both RAREs and CRs in space-ground communication scenarios. The receive signal-to-noise ratio (SNR) is required to be greater than $10$ dB, which is calculated by the link budget equation: ${\rm SNR}=\frac{P_{t}GL}{P_{N}}$. The antenna gain $G$ is set as 30 dBi, and the path loss $L$ scales as $d^{-2.3}$ according to the tyical space-ground channel model. The noise power, $P_{N}$, is proportional to the square of receiver sensitivity limit, $E^{2}_{\min}$. The frequency is fixed as 6 GHz by the electron transition $70D_{5/2}\leftrightarrow 71P_{3/2}$. Fig. 4 presents the coverage-expansion capability of a SQL-approaching RARE. When $P_{t}=5\times 10^{5}$ mW and the bandwidth is 100 kHz, the coverage range can be extended from the medium-earth-orbit ($d\approx 7000$ km) up to the geostationary orbit ($d\approx 36000$ km). Moreover, it is clear from Fig. 4 that a RARE can save a large amount of transmit power for fixed coverage. If the coverage range is fixed at the medium-earth-orbit ($d=$5000 km) with a bandwidth of 1 MHz, the transmit power can be reduced from $450$ W down to $40$ W. This reveals the appealing applications of RAREs in long-range communications, like the SAGIN.
Figure 4: Comparison of feasible communication distance between RAREs and CRs. LEO, MEO, and GEO refer to the heights of low-earth-orbit, medium-earth-orbit, and geostationary satellites.
III-D Bandwidth
In current RARE-related experiments, the terminology “bandwidth” is abused into two concepts. First, the term “operating bandwidth” refers to the range of frequency that a receiver can listen to. Second, the concept of “instantaneous bandwidth” characterizes the highest wireless communication rate that a receiver can support, which follows the standard definition of “bandwidth” in the communication community.
III-D1 Operating bandwidth
The broad operating bandwidth is one advantage of RAREs. Specifically, the operating bandwidth of conventional dipole antennas is limited to several GHz due to their wavelength-dependent sizes. For example, a wideband dipole antenna centered at 10 GHz typically has an operating bandwidth of no more than 2.5 GHz. In contrast, Rydberg atoms possess abundant energy levels, allowing electrons to interact with various radio frequencies via the transition between different energy levels. For instance, the electron transitions $56D_{5/2}\leftrightarrow 57P_{3/2}$ and $56D_{5/2}\leftrightarrow 52F_{3/2}$ can resonate with the radio frequencies 12.01 GHz and 116.05 GHz, respectively. It has been calculated that a single RARE is capable of sensing a broad spectrum, ranging from MHz to THz bands [4].
III-D2 Instantaneous bandwidth
The communication rate of a RARE is restricted by several factors, such as the response time for a EIT process to become steady. As reported by exciting studies, the typical instantaneous bandwidth of a RARE is around 10 MHz [4], comparable to the bandwidth of 4G LTE (i.e., 1$\sim$20 MHz). Some recent efforts are endeavored to increase the RARE’s bandwidth. For example, researchers have proposed a spatiotemporal multiplexing technique of probe lasers to decrease the EIT response time [12]. This technique expands the RARE’s bandwidth to 100 MHz, approaching the bandwidth of sub-6GHz systems specified by 5G standards (i.e., 20$\sim$100 MHz). To date, continuously improving RARE’s instantaneous bandwidth remains a hot topic.
IV Recent Progress in RARE Techniques
Recently, there have been studies to develop advanced RARE techniques to enhance the communication and sensing performance. This section provides an overview of the state-of-the-art designs.
IV-A Frequency-division-Multiplexed RAREs
The FDM technology plays a crucial role in 5G wideband systems for mitigating multi-path interference and efficient utilization of spectrum resources. The advancement of FDM-RARE is thus regarded as the key enabler of RARE-empowered wideband communications. The state-of-the-art FDM-RAREs can be separated into two categories: standard and cross-band FDM-RAREs.
IV-A1 Standard FDM-RARE
As illustrated in Fig. 5, a standard FDM-RARE aims to detect standard FDM signals, where all sub-carrier frequencies lie within a bandwidth of several MHz with a kHz-level frequency spacing. In this context, a uniform electron transition |$\ket{3}\leftrightarrow\ket{4}$ is sufficient to down-convert all sub-carrier signals into the Rabi frequency. The major challenge in facilitating standard FDM-RAREs is that the nonlinearity of Rabi frequency, $\Omega(t)$, w.r.t the EM wave destroys the orthogonality among different sub-carriers, resulting in significant intercarrier crosstalk. To address this issue, a deep learning model was trained in [10] to directly infer FDM symbols from the received probe-laser power. The experiment therein considered four subcarriers with a frequency spacing of 2 kHz in line-of-sight environment. It was shown that the learning approach can demodulate binary phase-shift keying symbols with over 99% accuracy.
IV-A2 Cross-band FDM-RARE
The cross-band FDM-RARE leverages the broad operating bandwidth of RAREs to serve as an integrated full-frequency communication platform. On this platform, signal frequencies can be distributed across distinct frequency bands, with band spacing ranging from GHz to THz, as presented in Fig. 5. To enable signal detection, the RARE needs to couple each band to a unique energy-level pair, denoted as |$\ket{3}\leftrightarrow\ket{n},n\in\{4,5,\cdots,N\}$. The electron transition between each energy-level pair down-converts the coupled band by its respective transition frequency $f_{\ket{3}\rightarrow\ket{n}}$, thereby simultaneously converting the entire cross-band RF signal to the baseband. Utilizing this principle, researchers have demonstrated the co-detection of cross-band signals over 5 separate bands (1.72, 12.11, 27.42, 65.11, and 115.75 GHz) with a 3-dB instantaneous bandwidth of 6 MHz, verifying RARE’s full-frequency sensing capability [6].
Figure 5: Standard and cross-band FDM-RAREs.
IV-B Atomic-MIMO Communications
As a core technology in 5G and Beyond, MIMO communications can greatly enhance the spectral efficiency through the diversity and multiplexing gains. This motivates the research on atomic-MIMO communications, which utilizes an array of RAREs for receiving wireless data to boost both the sensitivity and channel capacity of RAREs.
In the literature, a simpler atomic single-input-multiple-output (SIMO) system was developed in [13] for the first time to explore the diversity gain, wherein four RAREs were used to jointly recover AM symbols from a single-antenna user. Experiments validated that the receive SNR scales linearly with the number of RAREs, as in traditional systems. To further reap the multiplexing gain, the current authors have investigated the multi-user signal detection capability in atomic-MIMO systems [8], where the EIT-heterodyne scheme was adopted to retrieve Quadrature-Amplitude-Modulated (QAM) symbols. It was proved that atomic-MIMO exhibits a nonlinear phase-retrieval transmission model, based on which an expectation-maximization algorithm was designed to accurately recover multi-user signals, demonstrating the feasibility of atomic-MIMO communications. In a followup work, the optimal MIMO precoding has been designed to achieve the channel capacity of atomic-MIMO [14]. It made the assumption to the EIT-heterodyne scheme that the reference signal shares the same frequency with the data signal and is much stronger. It was shown that the optimal precoder should process the I/Q baseband symbols independently, and the channel capacity scales as $\min(N_{r}/2,N_{t})\log{\rm SNR}$ ) for $N_{r}$ atomic receive antennas and $N_{t}$ classic transmit antennas. The discount factor $\frac{1}{2}$ imposed on $N_{r}$ arises because the Rabi frequency, $\Omega(t)$, only captures the magnitude of EM waves.
IV-C RARE-empowered Wireless Sensing
Given that a RARE is capable of detecting extremely weak signals, it can be deployed in wireless sensing systems to refine the sensing accuracy by orders of magnitude. Specifically, one existing prototype senses a target’s vibration from both WiFi (2.4 GHz) and mmWave (28 GHz) signals [5]. Therein, the EIT-heterodyne scheme, as illustrated in Fig. 2, was used to recover the phase of an EM wave, ${\mathbf{E}}(t)$, and then retrieve the vibration information. Experiment results show that the RARE can increase the sensing accuracy by more than 10 times w.r.t conventional receivers implemented on the USRP, LimdSDR, and Intel 5300 WiFi card. In addition to vibration sensing, RARE can also be used to sense the angle-of-arrival (AoA) for the purpose of localization. By measuring the signal phases at two different locations via the EIT-heterodyne scheme, the AoA of the signal can be obtained with a precision of $2.1^{\circ}$ [4].
IV-D Quantum Many-body RAREs for External-noise Mitigation
A RARE is characterized by low internal noise, as previously discussed. However, it is still affected by external noises resulting from the inter-user and inter-cell interferences. This problem introduces an obstacle to the sensitivity advantage of RAREs. In recent years, several studies have been carried out to mitigate the external noise.
One intriguing approach leverages the quantum many-body effect of Rydberg atoms to counterintuitively convert the power of external noise into that of wireless signal, achieving a noise-enhanced RARE [15]. In the approach, a collection of Rydberg atoms are interacted with each other. This many-body interaction transforms the original linear Schrödinger equation into a nonlinear one, which is shown to exhibit a bistability behavior. The incoming wireless signal could stimulate the system to switch between the two stable states. Then, by introducing a certain amount of external noise into the receiver, the switching magnitude can be amplified such that the output SNR is drastically improved. Experiment results demonstrate that the noise-enhanced RARE can surpass the sensitivity of traditional RAREs by 6.6 dB in the presence of external noise. This work opens a new promising direction for advancing RARE: the detrimental impact of external noise can be turned to a beneficial effect through the nonlinear many-body interactions in Rydberg atoms.
V Future Directions
Despite the rapid development of RARE in recent years, its integration with practical communication systems remains at a nascent stage. Numerous wireless-communication related open problems should be addressed to unleash the full potential of RAREs. A few are discussed as follows.
V-A Generalized FDM-RARE Techniques
While initial solutions [10, 6] exist for both standard and cross-band FDM-RAREs, the systematic design of FDM-RARE technology, involving channel estimation, signal detection, and synchronization, remains an unaddressed and crucial challenge. For standard FDM-RAREs, generalizing the learning method in [10] to accommodate hundreds of subcarriers in a multi-path environment is difficult due to the unexplainability of neural networks. Regarding cross-band FDM-RAREs, it is noteworthy that the quantum system has evolved from a four-level system to an $N$-level system [6]. However, at the time of writing, a comprehensive transmission model for this new $N$-level system has yet to be developed, and the potential crosstalk between cross-band signals remains unexplored. To address the above issues, it is essential to start from first principles and systematically develop transmission methods tailored for FDM-RAREs by thoroughly investigating their nonlinear transmission models.
V-B Enhanced Atomic-MIMO Communications
Figure 6: Enhanced atomic-MIMO communications.
The research on atomic-MIMO communication is still at its early stage. Numerous signal processing techniques at both the transmitter (Tx) and receiver (Rx) sides can be investigated to improve the communication performance, as shown in Fig. 6.
First, consider the Rx side. Recall that a RARE might be vulnerable to external interference. Apart from the aforementioned quantum many-body method, the development of signal combining algorithms tailored for atomic-MIMO presents a viable approach to mitigating inter-user interference. Furthermore, using RAREs for precise estimation of far-field and near-field channels is crucial for effective atomic-MIMO communications, but remains an uncharted area. Lastly, investigating the integration of atomic-MIMO and FDM-RARE techniques could unlock RARE’s capability of full multi-user and multi-frequency signal detection.
Next, developing Tx-side signal processing methods optimized for RAREs can also significantly improve the communication performance. To this end, one could relax some restrictive assumptions made in the MIMO precoding study in [14] we discussed earlier and explore the general EIT-heterodyne scheme, where the reference signal can have a different frequency and comparable strength to the user signal. Within this framework, optimizing the joint power allocation among the unknown user signal, ${\mathbf{E}}(t)$, and known reference signal, ${\mathbf{E}}_{r}(t)$, designing precoding for ${\mathbf{E}}(t)$, and refining the waveform of ${\mathbf{E}}_{r}(t)$, are all potential approaches for capacity enhancement. Moreover, it is promising to introduce diverse beamforming technologies to cooperate with RAREs, such as fluid-antenna, reconfigurable-intelligent-surface, and continuous-aperture-array empowered beamforming.
V-C Atomic Integrated Communication and Sensing
The extreme sensitivity of RAREs to both communication and sensing signals can facilitate the exciting application of atomic Integrated Communication and Sensing (ISAC). To advance this direction, the signal waveform utilized in conventional ISAC systems, like the orthogonal-chirp-division-multiplexing signal, need to be adjusted to align with the nonlinear transmission model revealed by Rabi frequency. Additionally, it is promising to exploit the cross-band response capability of Rydberg atoms in atomic-ISAC systems. By exciting electrons to proper energy levels, it becomes feasible to simultaneously utilize sub-6GHz signals to ensure communication coverage and high-frequency signals to refine sensing accuracy.
V-D Advancing Quantum Many-body RAREs
A more thorough exploration of the quantum many-body effects of Rydberg atoms is essential for mitigating external noises. The preliminary study in [15] focused on periodic sinusoidal signals and white external noise, with the evaluation metric limited to SNR. Holistic theories are required to investigate the characteristics of this quantum many-body RARE under the influence of aperiodic communication signals (such as QAM signals and FDM signals), and thereby quantitatively analyze the reduction of practical color external noise (such as inter-user interference). In addition, analysis of the input-output mutual information should be carried out to unveil the communication gain brought by the many-body effect, rather than solely relying on the SNR. Last, it is intractable to derive an analytical transmission model from the nonlinear Schrödinger equation. Thereafter, it might be necessary to use artificial intelligence technologies, such as physics-informed neural networks, to develop signal detection and channel estimation algorithms for quantum many-body RAREs.
VI Conclusions
This article provides an overview of the fundamentals, advantages, progresses, and future directions for RARE-aided wireless communications. Representing a paradigm shift in receiver architecture, RAREs detect classical EM signals via quantum phenomena, contributing to improved sensitivity, enhanced coverage, and broadened operating bandwidth. RAREs can seamlessly integrate with both classical techniques, such as FDM and MIMO, as well as quantum techniques, like the many-body effect, to enhance communication performance. However, the advancement of RAREs requires more efforts from the communication community to systematically develop mature transmission models, theoretical analysis frameworks, and efficient communications algorithms, making RARE a practically useful technology for future wireless networks.
Biographies
Mingyao Cui is a PhD student at The University of Hong Kong, Hong Kong.
Qunsong Zeng is a research assistant professor with The University of Hong Kong, Hong Kong.
Kaibin Huang is a professor with The University of Hong Kong, Hong Kong.

### v2

@@
 Abstract
-The advancement of Rydberg Atomic REceiver (RARE) is driving a paradigm shift in electromagnetic (EM) wave measurement.
-RAREs utilize the electron transition phenomenon of highly-excited atoms to interact with EM waves, thereby enabling wireless signal detection.
-Operating at the quantum scale, such new receivers have the potential to breakthrough the sensitivity limit of classical receivers, sparking a revolution in physical-layer wireless communications.
-The objective of this paper is to offer insights into RARE-aided communication systems.
+Rydberg Atomic REceiver (RARE) is driving a paradigm shift in electromagnetic (EM) wave measurement by harnessing the electron transition phenomenon of Rydberg atoms.
+Operating at the quantum scale, such receivers have the potential to breakthrough the performance limit of classic receivers, sparking a revolution in physical-layer wireless communications.
+The objective of this paper is to offer insights into RARE-empowered communications.
 We first provide a comprehensive introduction to the fundamental principles of RAREs.
-Then, a thorough comparison between RAREs and classical receivers is conducted in terms of the antenna size, sensitivity, coverage, and bandwidth.
-Subsequently, we overview the state-of-the-art design in RARE-aided wireless communications, exploring the latest progresses in frequency-division multiplexing, multiple-input-multiple-output, wireless sensing, and quantum many-body techniques.
-Finally, we highlight several wireless-communication related open problems as important research directions.
+Then, a thorough comparison between RAREs and classic receivers is conducted in terms of the antenna size, sensitivity, bandwidth, and hardware cost.
+Subsequently, we overview the recent progresses in RARE-aided wireless communications, covering the frequency-division multiplexing, multiple-input-multiple-output, wireless sensing, and quantum many-body techniques.
+Moreover, the unique application of RARE in multiband sensing and communication is introduced.
+Finally, we conclude by providing promising research directions.
 Index Terms:
@@
 I Introduction
-The six-generation (6G) communication networks will leverage all available resources to support the explosive growth of global mobile services [1].
-To achieve seamless coverage, 6G will expand beyond terrestrial networks in 1G–5G to space–air–ground integrated networks (SAGINs) [1].
-Furthermore, the full spectrum, including sub-6GHz, millimeter-wave (mmWave), and terahertz (THz) bands, will be explored for attaining high-rate and multi-modal connections.
-These ambitious long-range and full-frequency communication visions require the next-generation receiver to exhibit exceptional sensitivity to electromagnetic (EM) waves across a wide spectrum.
-Since Hertz validated the existence of EM waves, classical receivers detect wireless signals primarily by using an antenna to transform incident EM waves into informative electric currents [2].
+The six-generation (6G) communication networks will leverage all available resources to support the explosive growth of global mobile services.
+To achieve seamless coverage, 6G will expand beyond terrestrial networks in 1G–5G to space–air–ground integrated networks.
+The full spectrum, including sub-6G, millimeter-wave (mmWave), and terahertz (THz) bands, will be explored for attaining high-rate and multi-modal connections.
+These ambitious visions require the next-generation receiver to exhibit exceptional sensitivity to electromagnetic (EM) waves across a wide spectrum.
+In current communication systems, receivers detect wireless signals by using an antenna to transform incident EM waves into electric currents [1].
 This mechanism, however, presents insurmountable limitations to physical-layer (PHY) wireless communications.
-Specifically, the front-end circuits of a receiver, such as amplifiers and mixers, introduce thermal noise into data currents.
-The noise power, -174 dBm/Hz, restricts the receiver’s sensitivity limit.
-In addition, an ideal dipole antenna typically has a half-wavelength size, which constrains its operating bandwidth to a narrow range around the center carrier frequency [1].
-To detect signals across the sub-6GHz, mmWave, and THz bands, the receiver has to equip an array of bulky antennas and font-end circuits, each dedicated to a specific band.
-Consequently, these challenges constrain the ultimate performance of wireless communications.
-Recently, at the intersection of quantum sensing and wireless communication, the Rydberg Atomic REceiver (RARE) has emerged as a radical solution to overcome these limitations [3].
+Firstly, the front-end circuits of a receiver, like amplifiers and mixers, introduce thermal noise into data currents, restricting the receiver’s sensitivity limit.
+In addition, an ideal dipole antenna typically has a half-wavelength size, which constrains its working bandwidth to a narrow range around the carrier frequency.
+To detect signals from sub-6G to THz, the receiver has to equip an array of bulky antennas and font-end circuits, each dedicated to a specific band.
+Consequently, these challenges constrain the ultimate performance of wireless communications [2].
+Recently, at the intersection of quantum sensing and wireless communication, Rydberg Atomic REceiver (RARE) has emerged as a radical solution to overcome these limitations [3].
 Rydberg atoms are highly-excited atoms with outer electrons far from the nucleus.
-Due to the large electric dipole moments, Rydberg atoms can strongly interact with incident EM waves, triggering electron transitions between resonate energy levels.
+Due to the large electric dipole moments, Rydberg atoms can strongly interact with incident EM waves, triggering electron transitions between energy levels.
 As a result, RAREs can accurately detect wireless signals by monitoring changes on the atoms’ quantum states resulting from electron transitions [2].
@@
 For instance, a RARE is immune to thermal noise as it involves no front-end circuits.
-Hence, its sensitivity limit can far exceed those of classical counterparts [5].
-Moreover, a single RARE is able to resonate with a large number of frequencies ranging from MHz to THz band because of the abundant energy levels in Rydberg atoms [6].
-These quantum advantages make RAREs a promising technology to revolutionize PHY communications in diverse applications, such as weak signal detection, coverage enhancement, power saving, and full-frequency sensing.
+Hence, its sensitivity limit can far exceed those of classic counterparts [4].
+Moreover, a single RARE is able to interact with a large number of frequencies ranging from Megahertz (MHz) to THz band because of the abundant energy levels in Rydberg atoms [5].
+These quantum advantages make RAREs a promising technology to revolutionize PHY communications in diverse applications, such as weak signal detection, coverage enhancement, and full-frequency sensing.
 The objective of this paper is to offer a beginner’s guide to the next-generation wireless communications aided by RAREs.
 To this end, we first comprehensively introduce the fundamental features of Rydberg atoms and the key modules of a RARE.
-Then, the performance of RAREs is compared with classical receivers in antenna size, sensitivity, coverage, and bandwidth.
-Subsequently, we overview the state-of-the-art RARE techniques for wireless communications.
-The latest progresses in four important directions are covered: frequency-division-multiplexing (FDM), multiple-input-multiple-output (MIMO), wireless sensing, and quantum many-body techniques.
-Last, we highlight several wireless-communication related open problems as important future directions to conduct research on RAREs.
+Then, the performance of RAREs is compared with classic receivers in antenna size, sensitivity, bandwidth, and hardware cost.
+Subsequently, we overview the state-of-the-art RARE techniques for wireless communications, covering frequency-division multiplexing (FDM), multiple-input-multiple-output (MIMO), wireless sensing, and quantum many-body techniques.
+Then, the potential of RARE in multiband sensing and communication services is revealed using experimental results.
+Last, we highlight promising future research directions on RARE from the PHY communication perspective.
 II Fundamentals of Rydberg Atomic Receivers
@@
 An atom consists of a positively charged nucleus orbited by negatively charged electrons.
-According to the Bohr model, electrons in these orbits have unique quantum states and quantized energy levels.
+Electrons in these orbits have unique quantum states and quantized energy levels.
 Each orbit is labeled by three quantum numbers $n\ell_{j}$, (e.g., $n\ell_{j}=5S_{1/2}$), including a principal quantum number $(n=1,2,3,\cdots)$, an azimuthal quantum number $(\ell=S,P,D,\cdots)$, and a total angular momentum quantum number $(j=\frac{1}{2},\frac{3}{2},\frac{5}{2},\cdots)$.
@@
 1).
-Then, Rydberg atom is defined as a highly-excited atom with one or more electrons that have a large principal quantum number, i.e., $n>20$ [4].
+Rydberg atom is defined as a highly-excited atom with one or more electrons that have a large principal quantum number, i.e., $n>20$.
 II-A2 Electron transition for preparing Rydberg atom
-Electrons can transit between energy levels by either absorbing or emitting photons from EM waves (or lasers) [5].
-This electron-transition process occurs when the photon’s frequency resonates (or nearly resonates) with the transition frequency between energy levels, which is defined as the ratio between the energy-level spacing and Plank constant.
+Electrons can transit between energy levels by absorbing photons from (or emitting photons to) EM waves [4].
+This electron-transition process occurs when the photon’s frequency resonates with the transition frequency between energy levels, which is defined as the ratio between the energy-level spacing and Plank constant.
 The process facilitates the creation of Rydberg atoms via the excitation of electrons to high energy levels.
@@
 To retrieve the wireless signal, it is essential to accurately readout the value of Rabi frequency.
-The standard measurement scheme is the EIT spectroscopy, developed based on the following two observations [9].
+The standard measurement scheme is the EIT spectroscopy, developed on the following observations [9].
   • The $(1,2)$-th element of the final quantum state, $\rho_{12}$, determines the rate at which Rydberg atoms absorb photons from the probe laser.
@@
 The interval of these two frequencies is shown to be linearly proportional to the Rabi frequency $\Omega(t)$ [3].
-These two observations suggest that by sweeping the probe-laser frequency around $f_{\ket{1}\rightarrow\ket{2}}$, two peaks can appear in the spectrum of the output probe laser, as depicted in Fig.
+These observations suggest that by sweeping the probe-laser frequency around $f_{\ket{1}\rightarrow\ket{2}}$, two peaks can appear in the spectrum of the output probe laser, as depicted in Fig.
 2.
@@
 2, both amplitude-modulated (AM) and frequency-modulated (FM) symbols can be illustrated directly using the maximum likelihood method due to the dependence of Rabi frequency on $\left|{\bm{\mu}^{T}{\mathbf{E}}(t)}\right|$ and $\delta$.
-For the detection of phase-modulated (PM) symbols, a EIT-heterodyne scheme was developed in [7], which originates from the holographic-phase sensing technique in optics imaging.
+For the detection of phase-modulated (PM) symbols, a heterodyne sensing scheme was developed in [6], which originates from the holographic-phase sensing technique in optics imaging.
 It introduces an auxiliary source to produce a known reference signal, ${\mathbf{E}}_{r}(t)$, which interferes with the unknown data signal, ${\mathbf{E}}(t)$, over the air.
@@
 Recent studies have also reported the detection of multi-frequency signals [10], multi-user signals [8], and target vibrations [4], all based on the principles we have covered and will be elaborated in Section IV.
-III Rydberg Atomic Receiver vs. Classical Receiver
+III Rydberg Atomic Receiver vs. Classic Receiver
 The RARE technique represents a paradigm shift in wireless signal detection from using electronic circuits to embracing quantum sensing.
@@
 III-A Antenna Size
-According to the Chu limit, a lossless dipole antenna in a classical receiver (CR) has a size of half-wavelength, $\frac{\lambda}{2}$.
+According to the Chu limit, a lossless dipole antenna in a classic receiver has a size of half-wavelength, $\frac{\lambda}{2}$.
 This wavelength-dependent size leads to bulky receivers at low frequencies, e.g., exceeding 10 cm when $f<1$ GHz.
@@
 The sensitivity of a receiver, $E_{\min}$, defined as the minimum detectable field strength per square root of averaging time (unit: $\rm V/m/\sqrt{Hz}$), is fundamentally limited by the internal noise generated by the receiver itself.
-For an ideal $\frac{\lambda}{2}$ dipole antenna of a CR, its internal noise stems from the thermal noise generated by front-end circuits, such as amplifiers and mixers.
+For an ideal $\frac{\lambda}{2}$ dipole antenna of a classic receiver, its internal noise stems from the thermal noise generated by front-end circuits, such as amplifiers and mixers.
 The corresponding sensitivity limit is expressed as $E_{\min}=\sqrt{\frac{P_{N}Z_{0}}{A_{e}}}$, where $P_{N}=-174{\rm dBm}/{\rm Hz}$ denotes the room-temperature thermal-noise power, $Z_{0}$ the free-space impedance, and $A_{e}$ the effective area.
-On the contrary, since a RARE contains no front-end circuits, its thermal noise is significantly weaker than a CR [5].
+On the contrary, since a RARE contains no front-end circuits, its thermal noise is significantly weaker than a classic receiver [4].
 Its major internal noise is the quantum shot noise (QSN) resulting from the randomness when observing quantum states [1].
@@
 The theoretical sensitivity limit of a RARE, also called Standard Quantum Limit (SQL), is calculated as $E_{\min}=\frac{\hbar}{|\bm{\mu}|\sqrt{N_{a}T_{r}}}$ [1].
-Figure 3: Sensitivity comparison between RAREs and CRs.
+Figure 3: Sensitivity comparison between RAREs and classic receivers.
 Fig.
-3 compares the sensitivity of RAREs with CRs.
+3 compares the sensitivity of RAREs with classic receivers.
 Representative values $T_{r}=225{\rm\mu s}$ and $N_{a}=5\times 10^{5}$ as in [3] are utilized.
@@
 It is anticipated that in the near future, RAREs will surpass the ideal $\frac{\lambda}{2}$ dipole antenna and approach SQL, with multi-wave-mixing emerging as a promising approach for performance improvement.
-III-C Long Coverage Range or Low Transmission Power
-Given the extreme sensitivity promised by SQL, RAREs can theoretically detect much weaker signals than CRs, making it possible to significantly expand communication coverage.
-To see it more clearly, we plot Fig.
-4 to illustrate the feasible communication distance, $d$, with respect to (w.r.t) the transmit power, $P_{t}$, for both RAREs and CRs in space-ground communication scenarios.
-The receive signal-to-noise ratio (SNR) is required to be greater than $10$ dB, which is calculated by the link budget equation: ${\rm SNR}=\frac{P_{t}GL}{P_{N}}$.
-The antenna gain $G$ is set as 30 dBi, and the path loss $L$ scales as $d^{-2.3}$ according to the tyical space-ground channel model.
-The noise power, $P_{N}$, is proportional to the square of receiver sensitivity limit, $E^{2}_{\min}$.
-The frequency is fixed as 6 GHz by the electron transition $70D_{5/2}\leftrightarrow 71P_{3/2}$.
-Fig.
-4 presents the coverage-expansion capability of a SQL-approaching RARE.
-When $P_{t}=5\times 10^{5}$ mW and the bandwidth is 100 kHz, the coverage range can be extended from the medium-earth-orbit ($d\approx 7000$ km) up to the geostationary orbit ($d\approx 36000$ km).
-Moreover, it is clear from Fig.
-4 that a RARE can save a large amount of transmit power for fixed coverage.
-If the coverage range is fixed at the medium-earth-orbit ($d=$5000 km) with a bandwidth of 1 MHz, the transmit power can be reduced from $450$ W down to $40$ W.
-This reveals the appealing applications of RAREs in long-range communications, like the SAGIN.
-Figure 4: Comparison of feasible communication distance between RAREs and CRs. LEO, MEO, and GEO refer to the heights of low-earth-orbit, medium-earth-orbit, and geostationary satellites.
-III-D Bandwidth
-In current RARE-related experiments, the terminology “bandwidth” is abused into two concepts.
-First, the term “operating bandwidth” refers to the range of frequency that a receiver can listen to.
-Second, the concept of “instantaneous bandwidth” characterizes the highest wireless communication rate that a receiver can support, which follows the standard definition of “bandwidth” in the communication community.
-III-D1 Operating bandwidth
-The broad operating bandwidth is one advantage of RAREs.
-Specifically, the operating bandwidth of conventional dipole antennas is limited to several GHz due to their wavelength-dependent sizes.
-For example, a wideband dipole antenna centered at 10 GHz typically has an operating bandwidth of no more than 2.5 GHz.
-In contrast, Rydberg atoms possess abundant energy levels, allowing electrons to interact with various radio frequencies via the transition between different energy levels.
-For instance, the electron transitions $56D_{5/2}\leftrightarrow 57P_{3/2}$ and $56D_{5/2}\leftrightarrow 52F_{3/2}$ can resonate with the radio frequencies 12.01 GHz and 116.05 GHz, respectively.
-It has been calculated that a single RARE is capable of sensing a broad spectrum, ranging from MHz to THz bands [4].
-III-D2 Instantaneous bandwidth
-The communication rate of a RARE is restricted by several factors, such as the response time for a EIT process to become steady.
-As reported by exciting studies, the typical instantaneous bandwidth of a RARE is around 10 MHz [4], comparable to the bandwidth of 4G LTE (i.e., 1$\sim$20 MHz).
-Some recent efforts are endeavored to increase the RARE’s bandwidth.
-For example, researchers have proposed a spatiotemporal multiplexing technique of probe lasers to decrease the EIT response time [12].
-This technique expands the RARE’s bandwidth to 100 MHz, approaching the bandwidth of sub-6GHz systems specified by 5G standards (i.e., 20$\sim$100 MHz).
+III-C Bandwidth
+In existing RARE studies, the term ”bandwidth” is often used to represent two distinct metrics.
+Firstly, it is utilized to define the range of frequencies that can trigger electron transitions of Rydberg atoms.
+This concept is commonly referred to as the “detectable frequency range” of a RARE in the literature.
+Secondly, it is also associated with the “instantaneous bandwidth”, which denotes the maximum data capacity that the receiver device can measure at any given moment.
+This term aligns with the standard definition of ”bandwidth” in the field of communication engineering.
+Typically, a RARE is characterized by a broad detectable frequency range but a narrow instantaneous bandwidth.
+III-C1 Detectable frequency range
+The detectable frequency range of a RARE can be significantly broader than that of a classic receiver.
+This is attributed to two key factors.
+Firstly, Rydberg atoms possess abundant energy levels.
+This property allows electrons to interact with various frequency bands, ranging from MHz to THz bands, via the transition between different energy levels [2].
+For instance, transitions $56D_{5/2}\leftrightarrow 57P_{3/2}$ and $56D_{5/2}\leftrightarrow 52F_{3/2}$ can resonate with the radio frequencies 12.01 GHz and 116.05 GHz, respectively.
+Furthermore, some quantum techniques, like the Zeeman effect and off-resonate ac Stark shift, can broaden the linewidth of electron transition.
+In this context, the detectable frequency range of a RARE can be further expanded to achieve continuous-frequency detection exceeding 1 GHz.
+III-C2 Instantaneous bandwidth
+Restricted by the long response time for an EIT process to become steady, the instantaneous bandwidth of a RARE is typically narrower than 10 MHz [7].
+This bandwidth is comparable to that of 4G LTE systems (1$\sim$20 MHz).
+Recent efforts are endeavored to increase the RARE’s bandwidth.
+For example, researchers have proposed a spatiotemporal multiplexing technique of probe lasers to decrease the EIT response time (see [139] of [2]).
+This technique expands the RARE’s bandwidth to 100 MHz, approaching the bandwidth of sub-6G systems specified by 5G standards (i.e., 20$\sim$100 MHz).
 To date, continuously improving RARE’s instantaneous bandwidth remains a hot topic.
-IV Recent Progress in RARE Techniques
+IV State-of-the-art RARE Techniques
 Recently, there have been studies to develop advanced RARE techniques to enhance the communication and sensing performance. This section provides an overview of the state-of-the-art designs.
-IV-A Frequency-division-Multiplexed RAREs
-The FDM technology plays a crucial role in 5G wideband systems for mitigating multi-path interference and efficient utilization of spectrum resources.
+IV-A Frequency-division Multiplexed RAREs
+The FDM technology plays a crucial role in 5G wideband systems for mitigating multi-path interference.
 The advancement of FDM-RARE is thus regarded as the key enabler of RARE-empowered wideband communications.
-The state-of-the-art FDM-RAREs can be separated into two categories: standard and cross-band FDM-RAREs.
+The state-of-the-art FDM-RAREs can be separated into two categories: standard and multi-band FDM-RAREs.
 IV-A1 Standard FDM-RARE
 As illustrated in Fig.
-5, a standard FDM-RARE aims to detect standard FDM signals, where all sub-carrier frequencies lie within a bandwidth of several MHz with a kHz-level frequency spacing.
-In this context, a uniform electron transition |$\ket{3}\leftrightarrow\ket{4}$ is sufficient to down-convert all sub-carrier signals into the Rabi frequency.
+4, a standard FDM-RARE aims to detect standard FDM signals, where all sub-carrier frequencies are compacted within a bandwidth of several MHz.
+In this context, a uniform electron transition |$\ket{3}\leftrightarrow\ket{4}$ is sufficient to down-convert all sub-carrier signals into the Rabi frequency of quantum state.
 The major challenge in facilitating standard FDM-RAREs is that the nonlinearity of Rabi frequency, $\Omega(t)$, w.r.t the EM wave destroys the orthogonality among different sub-carriers, resulting in significant intercarrier crosstalk.
@@
 It was shown that the learning approach can demodulate binary phase-shift keying symbols with over 99% accuracy.
-IV-A2 Cross-band FDM-RARE
-The cross-band FDM-RARE leverages the broad operating bandwidth of RAREs to serve as an integrated full-frequency communication platform.
+IV-A2 Multiband FDM-RARE
+The multiband FDM-RARE leverages the broad detectable frequency range of a RAREs to serve as an integrated full-frequency communication platform.
 On this platform, signal frequencies can be distributed across distinct frequency bands, with band spacing ranging from GHz to THz, as presented in Fig.
-5.
+4.
 To enable signal detection, the RARE needs to couple each band to a unique energy-level pair, denoted as |$\ket{3}\leftrightarrow\ket{n},n\in\{4,5,\cdots,N\}$.
-The electron transition between each energy-level pair down-converts the coupled band by its respective transition frequency $f_{\ket{3}\rightarrow\ket{n}}$, thereby simultaneously converting the entire cross-band RF signal to the baseband.
-Utilizing this principle, researchers have demonstrated the co-detection of cross-band signals over 5 separate bands (1.72, 12.11, 27.42, 65.11, and 115.75 GHz) with a 3-dB instantaneous bandwidth of 6 MHz, verifying RARE’s full-frequency sensing capability [6].
-Figure 5: Standard and cross-band FDM-RAREs.
+The electron transition between each energy-level pair down-converts the coupled band by its respective transition frequency $f_{\ket{3}\rightarrow\ket{n}}$, thereby simultaneously converting the entire multi-band RF signal to the baseband.
+Utilizing this principle, researchers have demonstrated the co-detection of 5 separate frequency bands (1.72, 12.11, 27.42, 65.11, and 115.75 GHz) with a 3-dB instantaneous bandwidth of 6 MHz, verifying RARE’s full-frequency sensing capability [5].
+Figure 4: Standard and multiband FDM-RAREs.
 IV-B Atomic-MIMO Communications
 As a core technology in 5G and beyond, MIMO communications can greatly enhance the spectral efficiency through the diversity and multiplexing gains.
-This motivates the research on atomic-MIMO communications, which utilizes an array of RAREs for receiving wireless data to boost both the sensitivity and channel capacity of RAREs.
+As a representative, the recent study in [12] has demonstrated a good performance with an 20 Gbps 1500 m transmission distance in free-space-optical communications using advanced MIMO diversity combining codes.
+The advancement of MIMO thus motivates the research on atomic-MIMO communications.
+Its key idea is to employ an array of RAREs, in place of traditional antenna arrays, for signal detection.
+In this context, both the high sensitivity of Rydberg atoms and the high spectral efficiency of MIMO can be leveraged.
 In the literature, a simpler atomic single-input-multiple-output (SIMO) system was developed in [13] for the first time to explore the diversity gain, wherein four RAREs were used to jointly recover AM symbols from a single-antenna user.
 Experiments validated that the receive SNR scales linearly with the number of RAREs, as in traditional systems.
-To further reap the multiplexing gain, the current authors have investigated the multi-user signal detection capability in atomic-MIMO systems [8], where the EIT-heterodyne scheme was adopted to retrieve Quadrature-Amplitude-Modulated (QAM) symbols.
-It was proved that atomic-MIMO exhibits a nonlinear phase-retrieval transmission model, based on which an expectation-maximization algorithm was designed to accurately recover multi-user signals, demonstrating the feasibility of atomic-MIMO communications.
-In a followup work, the optimal MIMO precoding has been designed to achieve the channel capacity of atomic-MIMO [14].
-It made the assumption to the EIT-heterodyne scheme that the reference signal shares the same frequency with the data signal and is much stronger.
-It was shown that the optimal precoder should process the I/Q baseband symbols independently, and the channel capacity scales as $\min(N_{r}/2,N_{t})\log{\rm SNR}$ ) for $N_{r}$ atomic receive antennas and $N_{t}$ classic transmit antennas.
-The discount factor $\frac{1}{2}$ imposed on $N_{r}$ arises because the Rabi frequency, $\Omega(t)$, only captures the magnitude of EM waves.
-IV-C RARE-empowered Wireless Sensing
+To further reap the multiplexing gain, [8] has investigated the multi-user signal detection capability of atomic-MIMO.
+It was proved that atomic-MIMO exhibits a nonlinear phase-retrieval transmission model.
+The nonlinear characteristic negates the efficacy of traditional zero-forcing and minimum-mean-square-error linear detectors, posing a challenge to atomic-MIMO signal detection.
+To this end, an Expectation-Maximization Gerchberg-Saxton algorithm was proposed to iteratively retrieve multi-user signals with a low complexity, demonstrating the feasibility of atomic-MIMO.
+IV-C RARE-empowered Quantum Wireless Sensing
 Given that a RARE is capable of detecting extremely weak signals, it can be deployed in wireless sensing systems to refine the sensing accuracy by orders of magnitude.
 Specifically, one existing prototype senses a target’s vibration from both WiFi (2.4 GHz) and mmWave (28 GHz) signals [4].
-Therein, the EIT-heterodyne scheme, as illustrated in Fig.
-2, was used to recover the phase of an EM wave, ${\mathbf{E}}(t)$, and then retrieve the vibration information.
-Experiment results show that the RARE can increase the sensing accuracy by more than 10 times w.r.t conventional receivers implemented on the USRP, LimdSDR, and Intel 5300 WiFi card.
-In addition to vibration sensing, RARE can also be used to sense the angle-of-arrival (AoA) for the purpose of localization.
-By measuring the signal phases at two different locations via the EIT-heterodyne scheme, the AoA of the signal can be obtained with a precision of $2.1^{\circ}$ [4].
+Therein, the heterodyne sensing scheme, as illustrated in Fig.
+2, was used to recover the phase of an EM wave and then retrieve the subtle-vibration of target.
+Experiment results show that the RARE can increase the sensing accuracy by more than 10 times as compared with conventional receivers implemented on the USRP, LimdSDR, and Intel 5300 WiFi card.
+In addition, RARE can also be utilized to sense angle-of-arrivals (AoA) and time-of-arrivals (ToA) for the purpose of localization.
+A self-heterodyne sensing technique was introduced in [14] to advance the current heterodyne sensing for high-resolution localization.
+This technique leverages the self-interference produced by the transmitted sensing signal as the reference signal, ensuring their waveforms to be identical.
+By doing so, it effectively overcomes the narrow instantaneous bandwidth limitation of RARE, enhancing the sensing bandwidth up to 100 MHz.
+Experiments show that a self-heterodyne RARE reduces the localization error by 20 dB compared with classic receivers.
 IV-D Quantum Many-body RAREs for External-noise Mitigation
@@
 Experiment results demonstrate that the noise-enhanced RARE can surpass the sensitivity of traditional RAREs by 6.6 dB in the presence of external noise.
-This work opens a new promising direction for advancing RARE: the detrimental impact of external noise can be turned to a beneficial effect through the nonlinear many-body interactions in Rydberg atoms.
-V Future Directions
-Despite the rapid development of RARE in recent years, its integration with practical communication systems remains at a nascent stage.
-Numerous wireless-communication related open problems should be addressed to unleash the full potential of RAREs.
-A few are discussed as follows.
-V-A Generalized FDM-RARE Techniques
-While initial solutions [10, 6] exist for both standard and cross-band FDM-RAREs, the systematic design of FDM-RARE technology, involving channel estimation, signal detection, and synchronization, remains an unaddressed and crucial challenge.
-For standard FDM-RAREs, generalizing the learning method in [10] to accommodate hundreds of subcarriers in a multi-path environment is difficult due to the unexplainability of neural networks.
-Regarding cross-band FDM-RAREs, it is noteworthy that the quantum system has evolved from a four-level system to an $N$-level system [6].
-However, at the time of writing, a comprehensive transmission model for this new $N$-level system has yet to be developed, and the potential crosstalk between cross-band signals remains unexplored.
-To address the above issues, it is essential to start from first principles and systematically develop transmission methods tailored for FDM-RAREs by thoroughly investigating their nonlinear transmission models.
-V-B Enhanced Atomic-MIMO Communications
-Figure 6: Enhanced atomic-MIMO communications.
-The research on atomic-MIMO communication is still at its early stage.
-Numerous signal processing techniques at both the transmitter (Tx) and receiver (Rx) sides can be investigated to improve the communication performance, as shown in Fig.
+This work opens a new promising direction for advancing RARE: the detrimental impact of external noise can be turned to a beneficial effect through the quantum many-body interactions.
+V RARE for Multiband Sensing and Communication (RARE-MSAC)
+Figure 5: A single RARE is used to simultaneously detect communication signals at low and medium frequency bands, and sensing signals at high frequency bands.
+We leverage the advancement of RARE to introduce a novel application in PHY wireless systems——the RARE-MSAC.
+The key idea is to exploit the multi-level electron transitions of Rydberg atoms for detecting multiband signals with diverse sensing and communication (S&C) requirements.
+As illustrated in Fig.
+5, a single RARE is employed.
+The atoms are prepared in an appropriate initial Rydberg state, |$\ket{3}$, where the transitions from |$\ket{3}$ to other Rydberg states, |$\ket{n},\forall n\geq 4$, can be coupled to distinct frequency bands.
+In this context, the RARE is capable of concurrently detecting communication signals at low ($<6{\rm GHz}$) or medium bands ($7{\rm GHz}-24{\rm GHz}$) to ensure coverage.
+Simultaneously, it can also capture sensing signals operating in the mmWave and THz bands to perform sensing tasks with varying granularity requirements, ranging from ToA estimation to heartbeat detection.
+It is amazing that all these S&C services are achieved by a single RARE device, which is nearly impossible for classic receivers due to the largely separated frequencies.
+Fig.
+6 compares the MSAC performance of the RARE and classic receivers.
+We consider a dual-band S&C system.
+A sub-6G device transmits communication signals operating at $3.213\>{\rm GHz}$ with an instantaneous bandwidth of $500\>{\rm kHz}$.
+Simultaneously, a mmWave device transmits sensing signals operating at $30.618\>{\rm GHz}$ with a bandwidth of $100\>{\rm kHz}$.
+Similar to [4], the sensing signal is reflected by a target for detecting its vibration .
+A single RARE is deployed to detect these dual-band signals for providing S&C services.
+The Rydberg states are configured as |$\ket{3}=60D_{5/2}$, |$\ket{4}=61P_{3/2}$, and |$\ket{5}=62P_{3/2}$, respectively.
+The electron transition $60D_{5/2}\rightarrow 61P_{3/2}$ has a dipole moment of $2.04\times 10^{-26}\>{\rm C\cdot m}$ and resonates with the $3.213\>{\rm GHz}$ communication signal, while the transition $60D_{5/2}\rightarrow 62P_{3/2}$ has a dipole moment of $6.24\times 10^{-27}\>{\rm C\cdot m}$ and resonates with the $30.618\>{\rm GHz}$ sensing part.
+The power of the probe laser and coupling laser are set as $120{\rm\mu W}$ and $34\>{\rm mW}$, respectively.
+For comparison, two classic receivers are also deployed to independently complete the aforementioned S&C tasks, denoted as “CR1” for the sub-6G communication service and “CR2” for the mmWave sensing service in Fig.
 6.
-First, consider the Rx side.
-Recall that a RARE might be vulnerable to external interference.
-Apart from the aforementioned quantum many-body method, the development of signal combining algorithms tailored for atomic-MIMO presents a viable approach to mitigating inter-user interference.
-Furthermore, using RAREs for precise estimation of far-field and near-field channels is crucial for effective atomic-MIMO communications, but remains an uncharted area.
-Lastly, investigating the integration of atomic-MIMO and FDM-RARE techniques could unlock RARE’s capability of full multi-user and multi-frequency signal detection.
-Next, developing Tx-side signal processing methods optimized for RAREs can also significantly improve the communication performance.
-To this end, one could relax some restrictive assumptions made in the MIMO precoding study in [14] we discussed earlier and explore the general EIT-heterodyne scheme, where the reference signal can have a different frequency and comparable strength to the user signal.
-Within this framework, optimizing the joint power allocation among the unknown user signal, ${\mathbf{E}}(t)$, and known reference signal, ${\mathbf{E}}_{r}(t)$, designing precoding for ${\mathbf{E}}(t)$, and refining the waveform of ${\mathbf{E}}_{r}(t)$, are all potential approaches for capacity enhancement.
-Moreover, it is promising to introduce diverse beamforming technologies to cooperate with RAREs, such as fluid-antenna, reconfigurable-intelligent-surface, and continuous-aperture-array empowered beamforming.
-V-C Atomic Integrated Communication and Sensing
-The extreme sensitivity of RAREs to both communication and sensing signals can facilitate the exciting application of atomic Integrated Communication and Sensing (ISAC).
-To advance this direction, the signal waveform utilized in conventional ISAC systems, like the orthogonal-chirp-division-multiplexing signal, need to be adjusted to align with the nonlinear transmission model revealed by Rabi frequency.
-Additionally, it is promising to exploit the cross-band response capability of Rydberg atoms in atomic-ISAC systems.
-By exciting electrons to proper energy levels, it becomes feasible to simultaneously utilize sub-6GHz signals to ensure communication coverage and high-frequency signals to refine sensing accuracy.
-V-D Advancing Quantum Many-body RAREs
-A more thorough exploration of the quantum many-body effects of Rydberg atoms is essential for mitigating external noises.
-The preliminary study in [15] focused on periodic sinusoidal signals and white external noise, with the evaluation metric limited to SNR.
-Holistic theories are required to investigate the characteristics of this quantum many-body RARE under the influence of aperiodic communication signals (such as QAM signals and FDM signals), and thereby quantitatively analyze the reduction of practical color external noise (such as inter-user interference).
-In addition, analysis of the input-output mutual information should be carried out to unveil the communication gain brought by the many-body effect, rather than solely relying on the SNR.
-Last, it is intractable to derive an analytical transmission model from the nonlinear Schrödinger equation.
-Thereafter, it might be necessary to use artificial intelligence technologies, such as physics-informed neural networks, to develop signal detection and channel estimation algorithms for quantum many-body RAREs.
-VI Conclusions
-This article provides an overview of the fundamentals, advantages, progresses, and future directions for RARE-aided wireless communications.
-Representing a paradigm shift in receiver architecture, RAREs detect classical EM signals via quantum phenomena, contributing to improved sensitivity, enhanced coverage, and broadened operating bandwidth.
-RAREs can seamlessly integrate with both classical techniques, such as FDM and MIMO, as well as quantum techniques, like the many-body effect, to enhance communication performance.
-However, the advancement of RAREs requires more efforts from the communication community to systematically develop mature transmission models, theoretical analysis frameworks, and efficient communications algorithms, making RARE a practically useful technology for future wireless networks.
+The transmission power for the two signals increases from $10\>{\rm dBm}$ to $40\>{\rm dBm}$.
+As observed from the experimental results of Fig.
+6, the spectrum efficiency in the communication service achieved by RARE is $\sim 6{\rm bps/Hz}$ higher than that of the classic receiver under Rayleigh fading channels.
+Additionally, the normalized mean square error (NMSE) of vibration sensing achieved by the RARE improves that achieved by the classic receiver by two orders of magnitude.
+Notably, these performance gains are accomplished by a single RARE without requiring additional RF circuits.
+Figure 6: Comparison of MSAC performance between a single RARE and two classic receivers, denoted as “CR 1” and “CR 2” respectively.
+VI Future Research Directions
+Despite the rapid development of RARE, its integration with practical communication systems remains at a nascent stage.
+Numerous physical-layer research problems deserve in-depth investigation to unleash the full potential of RAREs.
+  • More accurate transmission model.
+The current signal transmission model of RARE mainly considers the interaction between Rydberg atoms and EM waves.
+In practice, the Rydberg-Rydberg interaction can significantly influence the dynamics of quantum states, which should be thoroughly studied to make the signal transmission model of RARE more accurate.
+  • Joint transmitter and RARE optimization.
+As RARE changes the signal transmission model from linear to nonlinear, traditional signal processing methods become suboptimal.
+It is essential to develop new transmission techniques tailored for both the RARE and transmitter, including modulation, precoding, channel estimation, synchronization, interference mitigation, and so on.
+  • Physical layer security.
+RARE also holds promise in enhancing physical layer security.
+Taking advantage of the abundant energy levels, a RARE can flexibly switch the frequency band it is listening to.
+This capability enables the implementation of high-speed and ultra-broadband frequency hopping communication, which can effectively protect the privacy of wireless data.
+  • Satellite communication and low-altitude economy.
+Another promising application of RARE is to assist satellite communication and low-altitude economy.
+The high sensitivity of RARE not only enables the expansion of satellite signal coverage but also demonstrates the potential to enhance the precision of sensing and navigation for unmanned aerial vehicles.
+VII Conclusions
+This article provides an overview of RARE-aided wireless communications.
+Representing a paradigm shift in receiver architecture, RAREs detect EM signals via quantum phenomena, contributing to improved sensitivity and broadened detecable frequency range.
+RAREs can seamlessly integrate with both classic techniques, such as FDM and MIMO, as well as quantum techniques, like the many-body effect, to enhance communication performance.
+The potential of RARE in multiband S&C systems is revealed through experimental results.
+Finally, we have highlighted several promising future research directions on RARE.
 Biographies
-Mingyao Cui is a PhD student at The University of Hong Kong, Hong Kong.
+Mingyao Cui is a PhD candidate at The University of Hong Kong, Hong Kong.
 Qunsong Zeng is a research assistant professor with The University of Hong Kong, Hong Kong.

### v3

@@
 Abstract
-Rydberg Atomic REceiver (RARE) is driving a paradigm shift in electromagnetic (EM) wave measurement by harnessing the electron transition phenomenon of Rydberg atoms.
+Rydberg Atomic REceiver (RARE) is driving a paradigm shift in electromagnetic wave measurement by harnessing the electron transition phenomenon of Rydberg atoms.
 Operating at the quantum scale, such receivers have the potential to breakthrough the performance limit of classic receivers, sparking a revolution in physical-layer wireless communications.
-The objective of this paper is to offer insights into RARE-empowered communications.
+The objective of this paper is to offer insights into RARE-empowered communication systems.
 We first provide a comprehensive introduction to the fundamental principles of RAREs.
-Then, a thorough comparison between RAREs and classic receivers is conducted in terms of the antenna size, sensitivity, bandwidth, and hardware cost.
+Then, a thorough comparison between RAREs and classic receivers is conducted in terms of the antenna size, sensitivity, and bandwidth.
 Subsequently, we overview the recent progresses in RARE-aided wireless communications, covering the frequency-division multiplexing, multiple-input-multiple-output, wireless sensing, and quantum many-body techniques.
-Moreover, the unique application of RARE in multiband sensing and communication is introduced.
-Finally, we conclude by providing promising research directions.
+Moreover, the unique application of RARE in multi-band communication and sensing is unveiled, followed by numerical experiments to show its performance superiority over classic receivers by $\sim 2.4\>{\rm bps/Hz}$ in spectrum efficiency and $\sim 7.2\>{\rm dB}$ in sensing accuracy.
+Finally, we conclude this paper by providing promising research directions.
 Index Terms:
@@
 Firstly, the front-end circuits of a receiver, like amplifiers and mixers, introduce thermal noise into data currents, restricting the receiver’s sensitivity limit.
-In addition, an ideal dipole antenna typically has a half-wavelength size, which constrains its working bandwidth to a narrow range around the carrier frequency.
+In addition, an ideal dipole antenna typically has a half-wavelength size, which constrains its working bandwidth to a narrow range around the central frequency.
 To detect signals from sub-6G to THz, the receiver has to equip an array of bulky antennas and font-end circuits, each dedicated to a specific band.
-Consequently, these challenges constrain the ultimate performance of wireless communications [2].
+Consequently, these challenges constrain the performance upper-bound of wireless communications [2].
 Recently, at the intersection of quantum sensing and wireless communication, Rydberg Atomic REceiver (RARE) has emerged as a radical solution to overcome these limitations [3].
@@
 For instance, a RARE is immune to thermal noise as it involves no front-end circuits.
-Hence, its sensitivity limit can far exceed those of classic counterparts [4].
+Hence, its sensitivity limit can potentially exceed those of classic counterparts [4].
 Moreover, a single RARE is able to interact with a large number of frequencies ranging from Megahertz (MHz) to THz band because of the abundant energy levels in Rydberg atoms [5].
-These quantum advantages make RAREs a promising technology to revolutionize PHY communications in diverse applications, such as weak signal detection, coverage enhancement, and full-frequency sensing.
-The objective of this paper is to offer a beginner’s guide to the next-generation wireless communications aided by RAREs.
-To this end, we first comprehensively introduce the fundamental features of Rydberg atoms and the key modules of a RARE.
-Then, the performance of RAREs is compared with classic receivers in antenna size, sensitivity, bandwidth, and hardware cost.
-Subsequently, we overview the state-of-the-art RARE techniques for wireless communications, covering frequency-division multiplexing (FDM), multiple-input-multiple-output (MIMO), wireless sensing, and quantum many-body techniques.
-Then, the potential of RARE in multiband sensing and communication services is revealed using experimental results.
-Last, we highlight promising future research directions on RARE from the PHY communication perspective.
+These quantum advantages make RAREs a promising technology to revolutionize the next-generation wireless network.
+The objective of this paper is to provide a comprehensive guide for beginners to the RARE-empowered communication systems.
+To this end, we first introduce the fundamental characteristics of Rydberg atoms and the key components of a RARE.
+Next, the key indicators of RAREs are compared with classic receivers in terms of antenna size, sensitivity, and bandwidth.
+Subsequently, we review the state-of-the-art RARE techniques for wireless communications, covering frequency-division multiplexing (FDM), multiple-input-multiple-output (MIMO), quantum wireless sensing, and quantum many-body techniques.
+We then reveal a unique application of RAREs in multi-band communication and sensing systems—a single RARE can simultaneously detect communication signals at low frequencies to ensure coverage and sensing signals at high frequencies to enhance granularity.
+Numerical experiments confirm that RAREs outperform conventional counterparts, achieving a spectrum efficiency gain of $\sim 2.4\>{\rm bps/Hz}$ and a sensing accuracy improvement of $\sim 7.2\>{\rm dB}$.
+Last, we highlight promising future research directions on RARE from the PHY wireless communication perspective, including signal modeling, holographic atomic-MIMO, satellite communications, and related areas.
 II Fundamentals of Rydberg Atomic Receivers
@@
 First, the radius of an Rydberg atom is large due to its dependence on $n^{2}$ [2].
-For instance, the atomic radius grows from $150a_{0}$ to $10^{4}a_{0}$ as $n$ increases from $10$ to $100$, with $a_{0}=5.29\times 10^{-11}{\rm m}$ being the Bohr radius.
+For instance, the atomic radius grows from $150a_{0}$ to $10^{4}a_{0}$ as $n$ increases from 101010 to 100100100, with $a_{0}=5.29\times 10^{-11}{\rm m}$ being the Bohr radius.
 The substantial increase in radius endows the Rydberg atom with a large electric dipole moment, thus making it very sensitive to external EM wave.
 Second, the transition frequency between adjacent Rydberg states is abundant as the energy level spacing scales as $n^{-3}$ [3].
-For low energy levels ($n<20$), the transition frequency can reach the visible light frequency ($>100$ THz).
+For low energy levels ($n<20$), the transition frequency can reach the visible light frequency (>100>100> 100 THz).
 While for Rydberg states ($n>20$), the transition frequencies fall within the range of MHz to THz [4], exactly covering all the typical frequency bands used in PHY communications.
@@
 II-B2 EIT based quantum measurement
-Figure 2: EIT for measuring AM/FM/PM signals. For AM and PM detection, the detuning $\delta$ is zero, while for FM detection, the detuning $\delta$ is nonzero.
+Figure 2: EIT for measuring AM/FM/PM signals. For AM and PM detection, the detuning $\delta$ is zero, while for FM detection, $\delta$ is nonzero.
 To retrieve the wireless signal, it is essential to accurately readout the value of Rabi frequency.
 The standard measurement scheme is the EIT spectroscopy, developed on the following observations [9].
-  • The $(1,2)$-th element of the final quantum state, $\rho_{12}$, determines the rate at which Rydberg atoms absorb photons from the probe laser.
+  • The (1,2)(1,2)( 1 , 2 )-th element of the final quantum state, $\rho_{12}$, determines the rate at which Rydberg atoms absorb photons from the probe laser.
 The absorption loss of the probe-laser scales exponentially with $\rho_{12}$ [6].
@@
 So far, we have discussed the mechanism of detecting AM/FM/PM wireless signals using RAREs.
-Recent studies have also reported the detection of multi-frequency signals [10], multi-user signals [8], and target vibrations [4], all based on the principles we have covered and will be elaborated in Section IV.
+Recent studies have also reported the detection of multi-frequency signals [10], multi-user signals [8], and target sensing [4, 11], all based on the principles we have covered and will be elaborated in Section IV.
 III Rydberg Atomic Receiver vs. Classic Receiver
 The RARE technique represents a paradigm shift in wireless signal detection from using electronic circuits to embracing quantum sensing.
-This gives rise to a range of unique features for RARE-enabled PHY communications, which are introduced in this section.
+This gives rise to a range of unique features for RARE-enabled PHY wireless communications, which are introduced in this section.
 III-A Antenna Size
-According to the Chu limit, a lossless dipole antenna in a classic receiver has a size of half-wavelength, $\frac{\lambda}{2}$.
-This wavelength-dependent size leads to bulky receivers at low frequencies, e.g., exceeding 10 cm when $f<1$ GHz.
-In contrast, the vapor cells (i.e., antennas of RAREs) have wavelength-independent sizes, e.g., typically fixed as $1$ cm in diameter [6].
-Consequently, RAREs can have electrically small antennas, offering significant advantages especially for low-frequency applications.
+According to the Chu limit, a lossless and impedance-matched antenna in a conventional receiver requires a physical dimension proportional to the wavelength, e.g.
+$\frac{\lambda}{2}$, so as to efficiently convert EM waves into signals on a transmission line.
+At low frequencies, the size of an antenna element can reach hundreds of centimeters, for example exceeding $150\>{\rm cm}$ when $f<100\>{\rm MHz}$, leading to bulky receiver architectures.
+Conversely, a RARE does not absorb energy from the field, but rather leverages the atom-light interaction to perform nondestructive sensing.
+This mechanism enables the development of electrically small Rydberg antennas that are significantly smaller than the wavelength.
+Typically, the geometry size of a Rydberg antenna—comprising an ensemble of Rydberg atoms—is on the order of $1\>{\rm cm}\times 1\>{\rm mm}$, with the vapor cell length around $1\>{\rm cm}$ and the laser diameter approximately $1\>{\rm mm}$.
+This wavelength-independent size offers substantial advantages, especially for low-frequency applications.
 III-B Sensitivity Limit
@@
 Secondly, it is also associated with the “instantaneous bandwidth”, which denotes the maximum data capacity that the receiver device can measure at any given moment.
-This term aligns with the standard definition of ”bandwidth” in the field of communication engineering.
+This term aligns with the standard definition of ”bandwidth” in communication engineering.
 Typically, a RARE is characterized by a broad detectable frequency range but a narrow instantaneous bandwidth.
@@
 For example, researchers have proposed a spatiotemporal multiplexing technique of probe lasers to decrease the EIT response time (see [139] of [2]).
-This technique expands the RARE’s bandwidth to 100 MHz, approaching the bandwidth of sub-6G systems specified by 5G standards (i.e., 20$\sim$100 MHz).
+This technique expands the RARE’s bandwidth to 100 MHz, approaching that of sub-6G systems specified by 5G standards (i.e., 20$\sim$100 MHz).
 To date, continuously improving RARE’s instantaneous bandwidth remains a hot topic.
 IV State-of-the-art RARE Techniques
-Recently, there have been studies to develop advanced RARE techniques to enhance the communication and sensing performance. This section provides an overview of the state-of-the-art designs.
+There have been studies to develop advanced RARE techniques to enhance the communication and sensing performance.
+This section provides an overview of the most recent progresses of these state-of-the-art designs.
 IV-A Frequency-division Multiplexed RAREs
 The FDM technology plays a crucial role in 5G wideband systems for mitigating multi-path interference.
-The advancement of FDM-RARE is thus regarded as the key enabler of RARE-empowered wideband communications.
+The advancement of FDM-RARE is regarded as the key enabler of RARE-empowered wideband communications.
 The state-of-the-art FDM-RAREs can be separated into two categories: standard and multi-band FDM-RAREs.
@@
 It was shown that the learning approach can demodulate binary phase-shift keying symbols with over 99% accuracy.
-IV-A2 Multiband FDM-RARE
-The multiband FDM-RARE leverages the broad detectable frequency range of a RAREs to serve as an integrated full-frequency communication platform.
+IV-A2 Multi-band FDM-RARE
+The multi-band FDM-RARE leverages the broad detectable frequency range of a RAREs to serve as an integrated full-frequency communication platform.
 On this platform, signal frequencies can be distributed across distinct frequency bands, with band spacing ranging from GHz to THz, as presented in Fig.
@@
 Utilizing this principle, researchers have demonstrated the co-detection of 5 separate frequency bands (1.72, 12.11, 27.42, 65.11, and 115.75 GHz) with a 3-dB instantaneous bandwidth of 6 MHz, verifying RARE’s full-frequency sensing capability [5].
-Figure 4: Standard and multiband FDM-RAREs.
+Figure 4: Standard and multi-band FDM-RAREs.
 IV-B Atomic-MIMO Communications
 As a core technology in 5G and beyond, MIMO communications can greatly enhance the spectral efficiency through the diversity and multiplexing gains.
-As a representative, the recent study in [12] has demonstrated a good performance with an 20 Gbps 1500 m transmission distance in free-space-optical communications using advanced MIMO diversity combining codes.
+The recent study in [13] has demonstrated a 20 Gbps 1500 m transmission in free-space-optical communications using MIMO diversity combining codes.
 The advancement of MIMO thus motivates the research on atomic-MIMO communications.
@@
 By doing so, it effectively overcomes the narrow instantaneous bandwidth limitation of RARE, enhancing the sensing bandwidth up to 100 MHz.
-Experiments show that a self-heterodyne RARE reduces the localization error by 20 dB compared with classic receivers.
+Experiments show that a self-heterodyne RARE reduces the localization error by orders of magnitude compared with classic receivers.
 IV-D Quantum Many-body RAREs for External-noise Mitigation
@@
 Experiment results demonstrate that the noise-enhanced RARE can surpass the sensitivity of traditional RAREs by 6.6 dB in the presence of external noise.
-This work opens a new promising direction for advancing RARE: the detrimental impact of external noise can be turned to a beneficial effect through the quantum many-body interactions.
-V RARE for Multiband Sensing and Communication (RARE-MSAC)
-Figure 5: A single RARE is used to simultaneously detect communication signals at low and medium frequency bands, and sensing signals at high frequency bands.
-We leverage the advancement of RARE to introduce a novel application in PHY wireless systems——the RARE-MSAC.
-The key idea is to exploit the multi-level electron transitions of Rydberg atoms for detecting multiband signals with diverse sensing and communication (S&C) requirements.
+This work opens a new direction for advancing RARE: the detrimental impact of external noise can be turned to a beneficial effect through the quantum many-body interactions.
+V RARE for Multi-band Sensing and Communication (RARE-MSAC)
+Figure 5: A single RARE is used to simultaneously detect communication and sensing signals at different frequency bands.
+We leverage the advancement of RARE to introduce a novel application in wireless systems——the RARE-MSAC.
+The key idea is to exploit the multi-level electron transitions of Rydberg atoms for detecting multi-band signals with diverse sensing and communication (S&C) requirements.
 As illustrated in Fig.
@@
 Similar to [4], the sensing signal is reflected by a target for detecting its vibration.
-A single RARE is deployed to detect these dual-band signals for providing S&C services.
+A single RARE is deployed to detect these dual-band signals.
 The Rydberg states are configured as |$\ket{3}=60D_{5/2}$, |$\ket{4}=61P_{3/2}$, and |$\ket{5}=62P_{3/2}$, respectively.
 The electron transition $60D_{5/2}\rightarrow 61P_{3/2}$ has a dipole moment of $2.04\times 10^{-26}\>{\rm C\cdot m}$ and resonates with the $3.213\>{\rm GHz}$ communication signal, while the transition $60D_{5/2}\rightarrow 62P_{3/2}$ has a dipole moment of $6.24\times 10^{-27}\>{\rm C\cdot m}$ and resonates with the $30.618\>{\rm GHz}$ sensing part.
-The power of the probe laser and coupling laser are set as $120{\rm\mu W}$ and $34\>{\rm mW}$, respectively.
-For comparison, two classic receivers are also deployed to independently complete the aforementioned S&C tasks, denoted as “CR1” for the sub-6G communication service and “CR2” for the mmWave sensing service in Fig.
-6.
-The transmission power for the two signals increases from $10\>{\rm dBm}$ to $40\>{\rm dBm}$.
-As observed from the experimental results of Fig.
-6, the spectrum efficiency in the communication service achieved by RARE is $\sim 6{\rm bps/Hz}$ higher than that of the classic receiver under Rayleigh fading channels.
-Additionally, the normalized mean square error (NMSE) of vibration sensing achieved by the RARE improves that achieved by the classic receiver by two orders of magnitude.
+For comparison, two classic receivers are also deployed to independently complete the aforementioned S&C tasks, denoted as “CR1” for the communication service and “CR2” for the sensing.
+The transmission power increases from $-10\>{\rm dBm}$ to $10\>{\rm dBm}$.
+As observed in Fig.
+6, the spectrum efficiency in the communication service achieved by RARE is $\sim 2.4{\rm bps/Hz}$ higher than that of CR1 under Rayleigh fading channels.
+Additionally, the normalized mean square error (NMSE) of vibration sensing achieved by the RARE improves that achieved by CR2 by 7.2$\sim$ 7.2\>∼ 7.2dB.
 Notably, these performance gains are accomplished by a single RARE without requiring additional RF circuits.
-Figure 6: Comparison of MSAC performance between a single RARE and two classic receivers, denoted as “CR 1” and “CR 2” respectively.
+Figure 6: Comparison of MSAC performance between a single RARE and two classic receivers.
 VI Future Research Directions
@@
   • More accurate transmission model.
-The current signal transmission model of RARE mainly considers the interaction between Rydberg atoms and EM waves.
-In practice, the Rydberg-Rydberg interaction can significantly influence the dynamics of quantum states, which should be thoroughly studied to make the signal transmission model of RARE more accurate.
+The current signal transmission model of RARE primarily considers the interaction between Rydberg atoms and EM waves.
+In practice, Rydberg-Rydberg interactions can significantly influence the dynamics of quantum states.
+A comprehensive study of these interactions is essential to enhance the accuracy of the RARE signal transmission model.
   • Joint transmitter and RARE optimization.
-As RARE changes the signal transmission model from linear to nonlinear, traditional signal processing methods become suboptimal.
-It is essential to develop new transmission techniques tailored for both the RARE and transmitter, including modulation, precoding, channel estimation, synchronization, interference mitigation, and so on.
+As RARE modifies the signal transmission model from linear to nonlinear, traditional signal processing methods become suboptimal.
+It is essential to develop new transmission strategies tailored for both the RARE and transmitter, including modulation, precoding, channel estimation, synchronization, interference mitigation, and so on.
+  • Holographic atomic-MIMO communications.
+Conventionally, the performance of holographic MIMO—where massive antennas are deployed within a compact area—is limited by the mutual coupling between antenna elements.
+Fortunately, the holographic atomic-MIMO technique can overcome this limitation, as Rydberg atoms in different laser streams minimally influence each other.
+  • Data-driven RARE.
+To date, the quantum dynamics of Rydberg atoms under complicated EM environments, such as twisted radiowaves and broadband signals, remains largely unexplored due to the intricate Lindblad master equations.
+Data-drive methods, including physics-informed neural networks and foundation models, offer promising solutions for understanding and optimizing these RARE systems [10].
   • Physical layer security.
-RARE also holds promise in enhancing physical layer security.
-Taking advantage of the abundant energy levels, a RARE can flexibly switch the frequency band it is listening to.
-This capability enables the implementation of high-speed and ultra-broadband frequency hopping communication, which can effectively protect the privacy of wireless data.
+RARE holds significant potential for enhancing physical layer security.
+Leveraging abundant energy levels, RARE systems can flexibly switch their listening frequency bands.
+This capability enables high-speed and ultra-broadband frequency-hopping communication, effectively protecting wireless data privacy.
   • Satellite communication and low-altitude economy.
-Another promising application of RARE is to assist satellite communication and low-altitude economy.
+Another promising application of RARE is in supporting satellite communication and the low-altitude economy.
 The high sensitivity of RARE not only enables the expansion of satellite signal coverage but also demonstrates the potential to enhance the precision of sensing and navigation for unmanned aerial vehicles.
 VII Conclusions
-This article provides an overview of RARE-aided wireless communications.
-Representing a paradigm shift in receiver architecture, RAREs detect EM signals via quantum phenomena, contributing to improved sensitivity and broadened detecable frequency range.
+This article has provided an overview of RARE-aided wireless communications.
+Representing a paradigm shift in receiver architecture, RAREs detect EM signals via quantum phenomena, contributing to improved sensitivity and broadened detectable frequency range.
 RAREs can seamlessly integrate with both classic techniques, such as FDM and MIMO, as well as quantum techniques, like the many-body effect, to enhance communication performance.
-The potential of RARE in multiband S&C systems is revealed through experimental results.
+The potential of RARE in MSAC systems was revealed through experimental results.
 Finally, we have highlighted several promising future research directions on RARE.
+Acknowledgment
+This work was supported in part by the National Natural Science Foundation of China (No.
+624B2123), in part by the Research Grants Council of the Hong Kong Special Administrative Region, China under a fellowship award (HKU RFS2122-7S04), NSFC/RGC CRS (CRS_HKU702/24), the Areas of Excellence scheme grant (AoE/E-601/22-R), Collaborative Research Fund (C1009-22G), and the Grant 17212423, and in part by the Shenzhen-Hong Kong-Macau Technology Research Programme (Type C) (SGDX20230821091559018).
 Biographies